- HV card for R-HRS pre-shower need to be changed. Ramesh is working at it.
- Instructions to reset Q1 are not quite up to date: a broken link, we
had to turn the regulator off (for a while only) to be able to turn the
- A. Camsonne made a note about the status of the power switch hooked
to iocha11 (
- The white board calls for a raster size of 4*4 mm2 but in practice
3*3 mm2 is what is used. MCC is still using 6*6 in their units. Should
the board and the recovery procedure be updated?
- 8:00- started shift with production data taking in progress. iocha11
is not responding, it seems to us that this ioc is talking to the vdcs in
the spectrometer. As we are not taking data with these ones for now, we
treat this as a low priority item.
- 10:00 - MCC tried to get us up to 100 uA beam current, the beam got
very unstable. We have been running 80 since yesterday night and
apparently we will continue for a while.
- 11:30 - DAQ did not start 13974, kill coda according to instruction,
13975 started ok.
- 14:00 - Controlled access in the Hall to install shielding in front
of the Compton photon detector and power cycle the network switch is
hooked to iocha11 .
- 14:30- Left Q1 crashes. We follow instruction on the how to page to
reset it. The instructions are not quite up to date, Ramesh said he willl
- 15:00: going through the procedure for short beam down.
- 15:45: HV supply for the R-HRS pre-shower quits. Rameshis looking for
Still running slug 3 with so far 647k good pairs (run 13949 to 13970)
Parity run: k of good pair: k of accumulated good pairs
13971: 24 : 671 -- HRS: 25499, 5050
13972: 39 : 710
13973: 35 : 745
13974: 22 : 767
13975: daq hang up
13976: 39 : 806 -- HRS: 25502, 5052
13977: 37 : 843
13978: -- left arm only
Beam trip counter
10-11: 17 (got a little try at 90uA)